# Current address: ]College of William and Mary, Williamsburg, Virginia 23187-8795

### v1

@@ VII.4 Moments of the Deuteron Spin Structure Functions
Results for the deuteron moments $\overline{I}_{TT}^{d}$, $\overline{\Gamma}_{1}^{d}$, and $\overline{\gamma}_{0}^{d}$ are shown in Figs. 19, 20, and 21, respectively. The bar now indicates that, in addition to the elastic reaction, the deuteron electro-disintegration contribution is also excluded from the integrals. Therefore, they approximately correspond to the sum of the proton and neutron moments rather than to the deuteron nucleus (see Section II-E).
Figure 19: [Color online] EG4 results on the deuteron $\overline{\Gamma}_{1}^{d}(Q^{2})$. See Fig. 16 caption for details.
Figure 20: [Color online] EG4 results for the deuteron $\overline{I}_{TT}^{d}(Q^{2})$. The expected GDH value $\overline{I}_{TT}^{d}{(0)}=-1.574\pm 0.026$ is shown by the short horizontal line. Note that earlier CLAS data [85] and the phenomenological model predictions [48, 50] are not available for this integral. See Fig. 16 caption for more details.
Figure 21: [Color online] EG4 results on the deuteron $\overline{\gamma}_{0}^{d}(Q^{2})$. Note that phenomenological model predictions [48, 50] are not available for this integral. See Fig. 16 caption for more details.
Similar observations as for the proton measurements can be drawn: the EG4 data agree with the previous data from CLAS [85] (and from SLAC E143 [37], not shown in the figure). The measured $\overline{\Gamma}_{1}^{d}$ and $\overline{I}_{TT}^{d~{}\rm{exp}}(Q^{2})$ agree with the $\chi$EFT results for the sum of proton and neutron (multiplied with the $D$-state depolarization, see Section II-E) of Alarcón et al. [54], and of Bernard et al. [53] for the lowest $Q^{2}$ points. The models of Pasechnik et al. [50] and Burkert-Ioffe [47, 48] agree well with the $\overline{\Gamma}_{1}^{d}$ data, while the parameterization [85] is systematically slightly larger. The comparison between the MAID model and the data is similar to that of the proton for all three integrals.
@@ VII.5 Moments of the Neutron Spin Structure Functions
Moments of the neutron spin structure function can be extracted directly from those of the proton and neutron results, see Section II-E. Understanding the deuteron moments as “per nucleus” (rather than “per nucleon”), we rewrite Eqs. (36-38) as
@@
The resulting $\overline{\Gamma}_{1}^{n}(Q^{2})$ and $\overline{I}_{TT}^{n}(Q^{2})$ are shown in Figs. 22 and 23, respectively.
Figure 22: [Color online] EG4 results on the neutron $\overline{\Gamma}_{1}^{n}(Q^{2})$, compared with earlier results from EG1b (d,p) [85, 86] (black crossbars), JLab E97110 (neutron extracted from 3He) [65] (blue triangles), and E94010 (3He) [127] (green triangles). See Fig. 16 caption for more details.
Figure 23: [Color online] EG4 results on the neutron $\overline{I}_{TT}^{n}(Q^{2})$, compared with results from JLab E94010 (3He) [128] (green triangles) and E97110 (3He) [65] (blue triangles). The GDH value is shown as the short horizontal line at $\overline{I}^{n}_{TT}(0)=-0.915$. Note that earlier CLAS data [85] and the phenomenological model predictions [48, 50] are not available for this integral. See Fig. 16 caption for more details.
They agree with previous data from E143 [37], EG1a [129, 130] and EG1b [85, 86] where the neutron information was extracted from the combined deuteron and proton data, as well as those from E94010 [127] and E97110 [65] for which the neutron information was extracted from 3He data. There is also agreement with the phenomenological models of Burkert and Ioffe [47, 48] and Pasechnik et al. [50] (when available) and with the $\chi$EFT predictions from Bernard et al. [53] (up to $Q^{2}\simeq 0.08$ GeV2) and, to a lesser extent, with Alarcón et al. [54]. The MAID model prediction is systematically more positive than the data. Like in the case of $\overline{\gamma}_{0}^{n}$, the calculation of [124] (not shown in the figure) agrees in sign and behavior with the $\overline{\gamma}_{0}^{n}(Q^{2})$, up to a similar normalization factor as for $\overline{\gamma}_{0}^{p}(Q^{2})$.
@@
Figure 24: [Color online] EG4 results on the generalized spin polarizability $\overline{\gamma}_{0}^{n}$, compared with earlier results from JLab EG1b (d,p) [85, 86] (black crossbars), E97110 (3He) [65] (blue triangles), and E94010 (3He) [127] (green triangles). Note that phenomenological model predictions [48, 50] are not available for this integral. See Fig. 16 caption for more details.
@@ VII.6 Isospin Analysis on the Moments of Spin Structure Functions
Results on the proton and deuteron moments presented in the previous sections can be used to form various isospin combinations based on the description in Section II-E. The isoscalar ($p+n$) combination would be similar to the deuteron results, up to a factor $\approx 1/0.92$ from the deuteron $D$-state. For isovector ($p-n$) moments, the proton and deuteron results were combined as:
@@ VII.6.1 Results on the Bjorken Sum $\overline{\Gamma}_{1}^{p-n}(Q^{2})$
Figure 25: [Color online] The Bjorken sum $\overline{\Gamma}_{1}^{p-n}(Q^{2})$ from EG4 (solid red circles) and from the EG4 proton data combined with the JLab E97110 neutron data [65] (blue triangles). Also shown are earlier data from E94010/EG1a [43] (green triangles), CLAS EG1b [44] (black crossbars), and EG1dvcs [32] (brown open squares). The inner (outer) error bars on the data points give their statistical (total) uncertainties, while the systematic uncertainties of EG4 and EG4/E97110 are also given by the red and the blue bands, respectively. The $\chi$EFT predictions from Alarcón et al. [54] and Bernard et al. [53] (the latter without error band as given in Ref. [53]) are shown along with phenomenological model predictions from Burkert-Ioffe [47, 48], Pasechnik et al. [50], MAID [88], HLFQCD [52], and Gabdrakhmanov et al. [136].
@@
Both the EG4 and EG4/E97110 results on $\overline{\Gamma}_{1}^{p-n}$ are systematically above $\chi$EFT and model predictions, except that of HLFQCD [52] and [136], which agree with the data. To allow a quantitative comparison between the data and theory, we employed a fit of the form $bQ^{2}+cQ^{4}$ to the entire world data set, based on Eq. (32) [46], where $b$ is expected to be the GDH slope. Our results are shown in Table IV. The “$uncor$” uncertainty designates the point-to-point uncorrelated uncertainty. It is the quadratic sum of the statistical uncertainty and a fraction of the systematic uncertainty determined so that $\chi^{2}$/n.d.f = 1 for the best fit. The “$cor$” uncertainty is the correlated uncertainty estimated from the remaining fraction of the systematic uncertainty.
@@ VII.6.2 Isospin Study of $\overline{\gamma}_{0}(Q^{2})$
Figure 26: [Color online] EG4 results on the isovector (top) and isoscalar (bottom) combinations of the generalized longitudinal spin polarizability $\overline{\gamma}_{0}(Q^{2})$. The EG4-only data are shown by the red circles, compared with the earlier EG1b-only (black crossbars), the combination of EG4 $p$ and E97110 $n$(3He) data (blue triangles), and the combination of EG1b $p$ and E94010 $n$(3He) data (green triangles). The results are compared with $\chi$EFT predictions from Alarcón et al. [54] and Bernard et al., as well as the MAID model. The prediction by Bernard et al. [53] does not provide uncertainty bands for the proton-neutron difference; for illustration, we show a band whose width is the larger of the proton and the neutron uncertainties.
@@ Acknowledgements.
Supplementary material
IX Table for Results on $g_{1}^{p}$ and $A_{1}^{p}F_{1}^{p}$
@@
X Table for Results on $g_{1}^{D}$ and $A_{1}^{D}F_{1}^{D}$
@@
XI Table for Results on $g_{1}^{n}$ and $A_{1}^{n}F_{1}^{n}$
@@
XII Table for Results on Moments
@@
Table 8: Results for the proton moments $\overline{\Gamma}_{1}^{p}$, $\overline{I}_{TT}^{p}$ and $\overline{\gamma}_{0}^{p}$ in $10^{-4}$fm4, for the measured and the full ranges, along with their statistical and systematic uncertainties, as reported in [13]. For Figs. 4 and 5 of [13], the four lowest $Q^{2}$ bins were combined into two for visual clarity.
@@
Table 9: Results for the (per nucleus) deuteron moments $\overline{\Gamma}_{1}^{D}$, $\overline{I}_{TT}^{D}$ and $\overline{\gamma}_{0}^{D}$ in $10^{-4}$fm4, for the measured and the full ranges, along with their statistical and systematic uncertainties, as reported in [12].
@@
Table 10: Results for the neutron moments $\overline{\Gamma}_{1}^{n}$, $\overline{I}_{TT}^{n}$ and $\overline{\gamma}_{0}^{n}$ in $10^{-4}$fm4, extracted from the proton and the deuteron moments results as reported in [13, 12], along with their statistical and systematic uncertainties. The moments shown are for the full $x$ range.
@@
Table 11: Results on the Bjorken integral $\overline{\Gamma}_{1}^{p}-\overline{\Gamma}_{1}^{n}$: partial integral evaluated over the measured $x$-ranges, the near-full integral after adding an estimate of the unmeasured low-$x$ contributions down to $x=0.001$, and the uncertainties. The statistical uncertainty is from the measured data only, while the systematic uncertainty includes that from the low $x$ contribution.
@@
Table 12: Results on the isovector and isoscalar combinations $\overline{\gamma}_{0}^{p\mp n}$, in $10^{-4}$ fm4, along with their uncertainties.

### v2

@@ VII.4 Moments of the Deuteron Spin Structure Functions
Figure 19: [Color online] EG4 results on the deuteron $\overline{\Gamma}_{1}^{d}(Q^{2})$. The size of the systematic uncertainty (red colored band) varies from $0.0011$ to $0.0065$, while the relative size of the systematic to statistical uncertainty varies within $(44-98)\%$, from the lowest to the highest $Q^{2}$ bins. See Fig. 16 caption for more details.
Results for the deuteron moments $\overline{I}_{TT}^{d}$, $\overline{\Gamma}_{1}^{d}$, and $\overline{\gamma}_{0}^{d}$ are shown in Figs. 19, 20, and 21, respectively. The bar now indicates that, in addition to the elastic reaction, the deuteron electro-disintegration contribution is also excluded from the integrals. Therefore, they approximately correspond to the sum of the proton and neutron moments rather than to the deuteron nucleus (see Section II.5).
Figure 20: [Color online] EG4 results for the deuteron $\overline{I}_{TT}^{d}(Q^{2})$. Across all $Q^{2}$ bins, the size of the systematic uncertainty (red colored band) varies from $0.0198$ to $0.125$, while the relative size of the systematic to statistical uncertainty varies within $(42-83)\%$. The expected GDH value $\overline{I}_{TT}^{d}{(0)}=-1.574\pm 0.026$ is shown by the short horizontal line. Note that earlier CLAS data [102] and the phenomenological model predictions [54, 56] are not available for this integral. See Fig. 16 caption for more details.
Figure 21: [Color online] EG4 results on the deuteron $\overline{\gamma}_{0}^{d}(Q^{2})$. Across all $Q^{2}$ bins, the size of the systematic uncertainty (red colored band) varies from $0.0123$ to $0.339$, while the relative size of the systematic to statistical uncertainty varies within $(37-93)\%$. Note that phenomenological model predictions [54, 56] are not available for this integral. See Fig. 16 caption for more details.
Similar observations as for the proton measurements can be drawn: the EG4 data agree with the previous data from CLAS [102] (and from SLAC E143 [107], not shown in the figure). The measured $\overline{\Gamma}_{1}^{d}$ and $\overline{I}_{TT}^{d~{}\rm{exp}}(Q^{2})$ agree with the $\chi$EFT results for the sum of proton and neutron (multiplied with the $D$-state depolarization, see Section II.5) of Alarcón et al. [60], and of Bernard et al. [59] for the lowest $Q^{2}$ points. The models of Pasechnik et al. [56] and Burkert-Ioffe [53, 54] agree well with the $\overline{\Gamma}_{1}^{d}$ data, while the parameterization [102] is systematically slightly larger. The comparison between the MAID model and the data is similar to that of the proton for all three integrals.
@@ VII.5 Moments of the Neutron Spin Structure Functions
Moments of the neutron spin structure function can be extracted directly from those of the proton and neutron results, see Section II.5. Understanding the deuteron moments as “per nucleus” (rather than “per nucleon”), we rewrite Eqs. (36-38) as
@@
Figure 22: [Color online] EG4 results on the neutron $\overline{\Gamma}_{1}^{n}(Q^{2})$, compared with earlier results from EG1b (d,p) [102, 103] (black crossbars), JLab E97110 (neutron extracted from 3He) [71] (blue triangles), and E94010 (3He) [93] (green triangles). The size of the systematic uncertainty (red colored band) varies from $0.0013$ to $0.0071$, while the relative size of the systematic to statistical uncertainty varies from $(45-92)\%$, from the lowest to the highest $Q^{2}$ bins. See Fig. 16 caption for more details.
Figure 23: [Color online] EG4 results on the neutron $\overline{I}_{TT}^{n}(Q^{2})$, compared with results from JLab E94010 (3He) [92] (green triangles) and E97110 (3He) [71] (blue triangles). Across all $Q^{2}$ bins, the size of the systematic uncertainty varies from $0.0275$ to $0.141$, while the relative size of the systematic to statistical uncertainty varies within $(49-123)\%$. The GDH value is shown as the short horizontal line at $\overline{I}^{n}_{TT}(0)=-0.915$. Note that earlier CLAS data [102] and the phenomenological model predictions [54, 56] are not available for this integral. See Fig. 16 caption for more details.
The resulting $\overline{\Gamma}_{1}^{n}(Q^{2})$ and $\overline{I}_{TT}^{n}(Q^{2})$ are shown in Figs. 22 and 23, respectively. They agree with previous data from E143 [107], EG1a [94, 95] and EG1b [102, 103] where the neutron information was extracted from the combined deuteron and proton data, as well as those from E94010 [93] and E97110 [71] for which the neutron information was extracted from 3He data. There is also agreement with the phenomenological models of Burkert and Ioffe [53, 54] and Pasechnik et al. [56] (when available) and with the $\chi$EFT predictions from Bernard et al. [59] (up to $Q^{2}\simeq 0.08$ GeV2) and, to a lesser extent, with Alarcón et al. [60]. The MAID model prediction is systematically more positive than the data. Like in the case of $\overline{\gamma}_{0}^{n}$, the calculation of [154] (not shown in the figure) agrees in sign and behavior with the $\overline{\gamma}_{0}^{n}(Q^{2})$, up to a similar normalization factor as for $\overline{\gamma}_{0}^{p}(Q^{2})$.
@@
Figure 24: [Color online] EG4 results on the generalized spin polarizability $\overline{\gamma}_{0}^{n}$, compared with earlier results from JLab EG1b (d,p) [102, 103] (black crossbars), E97110 (3He) [71] (blue triangles), and E94010 (3He) [93] (green triangles). Across all $Q^{2}$ bins, the size of the systematic uncertainty varies from $0.0158$ to $0.397$, while the relative size of the systematic to statistical uncertainty varies from $(40-124)\%$. Note that phenomenological model predictions [54, 56] are not available for this integral. See Fig. 16 caption for more details.
@@ VII.6 Isospin Analysis on the Moments of Spin Structure Functions
Results on the proton and deuteron moments presented in the previous sections can be used to form various isospin combinations based on the description in Section II.5. The isoscalar ($p+n$) combination would be similar to the deuteron results, up to a factor $\approx 1/0.92$ from the deuteron $D$-state. For isovector ($p-n$) moments, the proton and deuteron results were combined as:
@@ VII.6.1 Results on the Bjorken Sum $\overline{\Gamma}_{1}^{p-n}(Q^{2})$
Figure 25: [Color online] The Bjorken sum $\overline{\Gamma}_{1}^{p-n}(Q^{2})$ from EG4 (solid red circles) and from the EG4 proton data combined with the JLab E97110 neutron data [71] (blue triangles). Also shown are earlier data from E94010/EG1a [42] (green triangles), CLAS EG1b [43] (black crossbars), and EG1dvcs [44] (brown open squares). The inner (outer) error bars on the data points give their statistical (total) uncertainties, while the systematic uncertainties of EG4 and EG4/E97110 are also given by the red and the blue bands, respectively. Across all $Q^{2}$ bins, the size of the systematic uncertainty of the EG4 results (red colored band) varies from $0.0017$ to $0.0115$, while the relative size of the systematic to statistical uncertainty varies from $(58-131)\%$. The $\chi$EFT predictions from Alarcón et al. [60] and Bernard et al. [59] (the latter without error band as given in Ref. [59]) are shown along with phenomenological model predictions from Burkert-Ioffe [53, 54], Pasechnik et al. [56], MAID [113], HLFQCD [58], and Gabdrakhmanov et al. [161].
@@
Both the EG4 and EG4/E97110 results on $\overline{\Gamma}_{1}^{p-n}$ are systematically above $\chi$EFT and model predictions, except that of HLFQCD [58] and [161], which agree with the data. To allow a quantitative comparison between the data and theory, we employed a fit of the form $bQ^{2}+cQ^{4}$ to the entire world data set, based on Eq. (32[52], where $b$ is expected to be the GDH slope. Our results are shown in Table 4. The “$uncor$” uncertainty designates the point-to-point uncorrelated uncertainty. It is the quadratic sum of the statistical uncertainty and a fraction of the systematic uncertainty determined so that $\chi^{2}$/n.d.f = 1 for the best fit. The “$cor$” uncertainty is the correlated uncertainty estimated from the remaining fraction of the systematic uncertainty.
@@ VII.6.2 Isospin Study of $\overline{\gamma}_{0}(Q^{2})$
Figure 26: [Color online] EG4 results on the isovector (top) and isoscalar (bottom) combinations of the generalized longitudinal spin polarizability $\overline{\gamma}_{0}(Q^{2})$. The EG4-only data are shown by the red circles, compared with the earlier EG1b-only (black crossbars), the combination of EG4 $p$ and E97110 $n$(3He) data (blue triangles), and the combination of EG1b $p$ and E94010 $n$(3He) data (green triangles). For $\overline{\gamma}_{0}^{p-n}(Q^{2})$, the size of the systematic uncertainty varies from $0.026$ to $0.522$, while the relative size of the systematic to statistical uncertainty varies from $(58-216)\%$. For $\overline{\gamma}_{0}^{p+n}(Q^{2})$, the size of the systematic uncertainty varies from $0.013$ to $0.367$, while the relative size of the systematic to statistical uncertainty varies from $(36-93)\%$. The results are compared with $\chi$EFT predictions from Alarcón et al. [60] and Bernard et al., as well as the MAID model. The prediction by Bernard et al. [59] does not provide uncertainty bands for the proton-neutron difference; for illustration, we show a band whose width is the larger of the proton and the neutron uncertainties.
@@ Acknowledgements.
Appendix A Table for Results on $g_{1}^{p}$ and $A_{1}^{p}F_{1}^{p}$
Results for $g_{1}^{p}$ and $A_{1}^{p}F_{1}^{p}$ as shown in Figs. 10 and 12, respectively.
@@ Appendix A Table for Results on $g_{1}^{p}$ and $A_{1}^{p}F_{1}^{p}$
Appendix B Table for Results on $g_{1}^{D}$ and $A_{1}^{D}F_{1}^{D}$
Results for $g_{1}^{d}$ and $A_{1}^{d}F_{1}^{d}$ as shown in Figs. 11 and 13, respectively.
@@ Appendix B Table for Results on $g_{1}^{D}$ and $A_{1}^{D}F_{1}^{D}$
Appendix C Table for Results on $g_{1}^{n}$
Results for $g_{1}^{n}$ as shown in Fig. 14.
@@ Appendix C Table for Results on $g_{1}^{n}$
Appendix D Table for Results on Moments
@@
Table 8: Results for the proton moments $\overline{\Gamma}_{1}^{p}$, $\overline{I}_{TT}^{p}$ and $\overline{\gamma}_{0}^{p}$ in $10^{-4}$fm4, for the measured and the full ranges, along with their statistical and systematic uncertainties, as shown in Figs.14-16 and reported in [22]. For Figs. 4 and 5 of [22], the four lowest $Q^{2}$ bins were combined into two for visual clarity.
@@
Table 9: Results for the (per nucleus) deuteron moments $\overline{\Gamma}_{1}^{D}$, $\overline{I}_{TT}^{D}$ and $\overline{\gamma}_{0}^{D}$ in $10^{-4}$fm4, for the measured and the full ranges, along with their statistical and systematic uncertainties, as shown in Figs. 19-21 and reported in [21].
@@
Table 10: Results for the neutron moments $\overline{\Gamma}_{1}^{n}$, $\overline{I}_{TT}^{n}$ and $\overline{\gamma}_{0}^{n}$ in $10^{-4}$fm4, extracted from the proton and the deuteron moments results as shown in Figs. 22-24 and reported in [22, 21], along with their statistical and systematic uncertainties. The moments shown are for the full $x$ range.
@@
Table 11: Results on the Bjorken integral $\overline{\Gamma}_{1}^{p}-\overline{\Gamma}_{1}^{n}$ as shown in Fig. 25: partial integral evaluated over the measured $x$-ranges, the near-full integral after adding an estimate of the unmeasured low-$x$ contributions down to $x=0.001$, and the uncertainties. The statistical uncertainty is from the measured data only, while the systematic uncertainty includes that from the low $x$ contribution.
@@
Table 12: Results on the isovector and isoscalar combinations $\overline{\gamma}_{0}^{p\mp n}$, in $10^{-4}$ fm4, along with their uncertainties, as shown in Fig. 26.